OpenStudy (anonymous):

Prove the following identity: (sin x)(tan x cos x - cot x cos x)=1-2cos^2 x

OpenStudy (anonymous):

\[tanx = \frac{ sinx }{ cosx }\] \[cotx = \frac{ cosx }{ sinx }\] Knowing this than you can substitute to get the following. \[(sinx)\frac{ sinx }{ cosx }(cosx) - \frac{ cosx }{ sinx }(cosx)\] now you can simplify to get..\[\sin ^{2}x - \frac{ \cos ^{2} x}{ sinx }\] can you take it from there knowing that \[\sin ^{2}x + \cos ^{2}x = 1 \]

OpenStudy (anonymous):

thank you